Circular No. 5705

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505

MARSDEN or GREEN@CFA.BITNET MARSDEN or GREEN@CFAPS2.SPAN

1993 BW_{3}

R. H. McNaught, Bugaldie, N.S.W., reports his discovery of a fastmoving asteroidal object on a U.K. Schmidt plate taken by M. J. Drinkwater. Follow-up observations with the Uppsala Southern Schmidt and all measures are by McNaught:

1993 UT  α_{2000}  δ_{2000}  V
Jan. 30.61345  10^{h}54^{m}00.^{s}02  −0^{o}45′06′′.5  17
30.65512  10 53 56.37  −0 45 16.9
Feb. 2.67832  10 49 04.95  −0 57 49.4
2.69360  10 49 03.32  −0 57 52.8
2.70549  10 49 02.07  −0 57 55.7
2.73187  10 48 59.58  −0 58 01.4
3.70938  10 47 22.55  −1 01 34.5

Orbital elements by the undersigned from six of the above observations:

T = 1992 Aug. 15.750 TT  ω = 76.^{o}409
e = 0.53726  Ω = 318.663 }2000.0
q = 1.01694 AU  i = 21.180

a = 2.19762 AU n^{o} = 0.302534 P = 3.26 years

1993 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
Jan. 23 11^{h}05.^{m}02 − 0^{o}03.^{′}3 1.139 1.965 135.^{o}5 20.^{o}6 17.3
Feb. 2 10 50.19 − 0 55.3 1.128 2.033 148.1 14.8 17.1
12 10 33.09 − 1 22.4 1.143 2.099 160.4  9.1 17.0
22 10 15.68 − 1 27.2 1.185 2.163 168.2  5.4 17.0
Mar. 4 9 59.99 − 1 16.0 1.257 2.226 163.3  7.4 17.3
14 9 47.48 − 0 56.8 1.357 2.286 152.6 11.5 17.7
24 9 38.78 − 0 36.5 1.481 2.345 141.5 15.3 18.0

PERIODIC COMET SCHAUMASSE (1992x) Total visual magnitude estimates: Jan. 16.78 UT, 11.6 (A. Renou,

Blaison, France, 0.25-m reflector); 19.12, 10.5 (C. Spratt, Victoria, B.C., 0.20-m reflector); 23.02, 9.8 (J. Carvajal, Madrid, Spain, 0.3-m reflector); 26.09, 8.9 (J. E. Bortle, Stormville, NY, 20 × 80 binoculars); 28.21, 8.8 (C. S. Morris, Whitaker Peak, CA, 20 × 80 binoculars).

1993 February 5 Gareth V. Williams